# Operators with small Kreiss constants

### v1

@@ Abstract.
We also study operators that satisfy a variant of the above Kreiss condition where $K$ is replaced by $1+\varepsilon(|z|)$, where the positive continuous function $\varepsilon(|z|)$ tends to 0 as $|z|\to 1^{+}.$ We show that, if the spectum of $T$ touches the unit circle only at a single point and the resolvent of $T$ satisfies a growth restriction along the unit circle, it is possible to choose $\varepsilon$ so that this Kreiss-type condition guarantees similarity to a contraction. At the core of our proof lies a positivity argument involving the double-layer potential operator. Counterexamples related to less restrictive choices of $\varepsilon$ are also provided.
@@ 1.1. The Kreiss condition
Kreiss [Kreiss62] proved that every matrix that satisfies the above condition is power-bounded. More specifically, for every $N\times N$ matrix $A$ with Kreiss constant $K$, there exists $M(N,K)>0$ with
@@
In applications, however, the dimension $N$ may be very large (see e.g. [Dorsselaer93]), so it is of interest to look for sharp constants $M(N,K)$.
Indeed, in the years after Kreiss published his result, various authors made successive improvements that reduced the value of $M(N,K)$ (see [pseudospectrabook, Section 18] for the full history and references). Tadmor [Tadmor81] was the first to show that $M(N,K)\leq CKN$ for some numerical constant $C>0$. Finally, Spijker [Spijkersharp], making crucial use of a Bernstein-type inequality for rational functions, proved that $M(N,K)\leq eKN$ and so, if $A$ is any $N\times N$ matrix with Kreiss constant $K,$ then
@@
LeVeque and Trefethen [LeVequeTref] showed that (1.2) is sharp in the sense that, for every $N\geq 1,$ there exists an $N\times N$ matrix $A_{N}$ with Kreiss constant $K_{N}$ such that
@@
By (1.2), we have $P(N,K)\leq eKN.$ What about lower bounds? The first results of this type are due to McCarthy-Schwartz [McCarthySchwartz] and van Dorsselaer-Kraaijevanger-Spijker [Dorsselaer93], who established lower bounds for $P(N,K)$ that were logarithmic in $N$. Later, Spijker, Tracogna and Welfert [STW]) almost bridged the gap with (1.2) by showing that, given $\varepsilon>0,$ there exists $K=K(\varepsilon)>\pi+1$ and $C=C(\varepsilon)>0$ such that
@@
Their construction was subsequently refined by Nikolski [Nikolski], who showed that, for any $K>\pi+1$ and $N\geq 1,$
@@
More recently, the search for non-power-bounded operators that satisfy various Kreiss- and Cesàro-type conditions has focused on weighted shifts; see e.g. [AbsolCesaromuller, KreissanduniKreiss, CohenCunyEisnerLin].
@@ 1.2. Small Kreiss constants
Unfortunately, all of the above-mentioned papers fall short of providing lower bounds for $P(N,K)$ for $K$ close to $1.$ In fact, it appears that the only time non-power-bounded operators with Kreiss-type constants close to $1$ were considered in the literature was in a paper of McCarthy [McCarthy71lost] (see also the remark in [Nikolski, p. 374]), who proved that, for any $\varepsilon>0,$ there exists $C=C(\varepsilon)$ with
@@
    Let $\alpha\in(0,\frac{1}{4})$. There exist $M=M(\alpha)>0$ and $\varepsilon_{0}\in(0,1]$ such that, whenever $0<\varepsilon<\varepsilon_{0}$ and $N^{\varepsilon}\geq M$, one has
@@ 1.3. The similarity problem
We now move from matrices to infinite-dimensional linear operators. Let $H$ be a Hilbert space and let $T\in\mathcal{B}(H)$ with $\sigma(T)\subset\overline{\mathbb{D}}$. The Kreiss condition
@@
(where $\nu>0$) no longer implies power-boundedness in this setting; classical counterexamples are contained in [Foguelcounterex, HalmosFoguel]. On the other hand, if $C=1$, then $T$ is a $\rho$-contraction for some $\rho>1$, and so it is similar to a contraction (and thus power-bounded), no matter the value of $\nu$ (see e.g. [Foiasbook]).
In the second half of our paper, we examine several conditions on $T$ that lie in-between the above $C=1$ and $C>1$ regimes, and determine whether they imply similarity to a contraction. We point out that similarity to a contraction (and, more generally, $K$-spectral estimates) has already been linked to resolvent conditions several times in the literature, see e.g. [BouabdLeMerdy, CassiergeneralizedToepl, deLauben, TestsforKsp, Gamal]. For more references and recent developments concerning the similarity problem, especially in relation to operator semigroups, see [mazatomilov]. One of the conditions we are interested in, naturally motivated by the consideration of “small” Kreiss constants, is
@@
We also consider operators $T$ such that $\lim_{n\to\infty}\|T^{n}\|=1$. Notice that if $\|T^{m}\|\leq 1$ for some $m\in\mathbb{N}$, then $T$ is similar to a contraction, see  [[Halmos-ten], discussion of Problem 6]. It is easy to see that the condition $\lim_{n\to\infty}\|T^{n}\|=1$ implies (1.9) for a certain function $\varepsilon(x)$ vanishing at $0^{+}$. By adapting Pisier’s example of an operator that is polynomially bounded but not similar to a contraction [Pisier-result], we show
@@
Let $T\in\mathcal{B}(H)$ with $\sigma(T)\subset\mathbb{D}\cup\{1\}$, with $\gamma:[-\delta_{-},\delta_{+}]\to\mathbb{C}$ a v-type curve. Assume that $T$ satisfies the Kreiss condition and, moreover, $\varepsilon:(0,\nu_{0}]\to(0,\infty)$ is a continuous increasing function such that $\varepsilon(0^{+})=0$ and the resolvent estimate (1.9) holds for all $\lambda$ in the set determined by $\gamma$ (excluding $\lambda=1$). Set $r(\theta)=|\gamma(\theta)|$. If
@@
for all $\theta$ (see e.g. [schwenntadritt, Section 1.2]). Thus, in view of Theorem 1.5, we immediately obtain
@@
We note that Ritt operators, while power-bounded, are not, in general, similar to contractions [LMsimprob, Proposition 5.2].
@@ 1.4. Outline
Our paper is organized as follows: in Section 2, we briefly go over the definitions and the relations between those Kreiss- and Cesáro-type conditions that are relevant to this work. We also record an inequality from [AbsolCesaromuller] that will be very useful in the proof of Theorem 1.2. Section 3 contains the proofs of Theorem 1.2 and Theorem 3.3, and, in addition, a lower bound corresponding to the strong Kreiss condition (Remark 3.7). Theorem 1.5 is proved in Section 4. Section 5 contains the proofs of Theorems 5.1 and 1.3. Section 6 concludes with some open questions related to our investigations.
@@ 2. Kreiss- and Cesàro-type conditions
Clearly, power-boundedness implies all of the above conditions. Absolute Cesàro boundedness and the strong Kreiss condition are both strictly stronger than the uniform Kreiss condition, see [AbsolCesaromuller, GomilkoZemanek, Mont], while the two are independent of one another [CohenCunyEisnerLin]. Further, the uniform Kreiss condition is equivalent [Mont] to the existence of a constant $C>0$ such that, for all $n\geq 1$ and $|\lambda|=1,$
@@
and is thus stronger than Cesáro boundedness (and also, trivially, stronger than the Kreiss condition). The Kreiss condition and Cesàro boundedness are independent [SuciuZem]. Finally, while the strong Kreiss condition does not imply power-boundedness [CohenCunyEisnerLin], we observe that the similar-looking condition
@@
is, in fact, equivalent to power-boundedness [Gibson, Packel].
@@
An interesting example of an operator that is absolutely Cesáro but not power bounded is given in [[AbsolCesaromuller], Theorem 2.1] (we only record the $\ell^{2}$-version here):
@@ 3.1. Lower bounds for the absolute Cesàro condition
$P_{AC}(M,1+\varepsilon)\geq 2^{-m}C(\varepsilon,m)\log^{m}(M),$ (3.14)
for all $M\geq m^{2}2^{2m}$. Indeed, let $M$ satisfy $M\geq m^{2}2^{2m},$ with $N^{\prime}$ the largest integer such that $N^{\prime}\leq M$ and $N^{\prime}$ is a multiple of $2^{m}$. Since $P_{AC}(M,K)$ is increasing with respect to $M$ (assuming $K$ fixed), (3.13) allows us to write
$P_{AC}(M,1+\varepsilon)$ $\geq P_{AC}(N^{\prime},1+\varepsilon)$
@@
$\geq C(\varepsilon,m)\log^{m}(M/2)$
$\geq 2^{-m}C(\varepsilon,m)\log^{m}(M),$
@@
Set $m=\lfloor\frac{v}{2\log v}\rfloor$, where $v=\log(\varepsilon\log N)$. This choice of $m$ is, in a sense, close to the value that maximizes this last lower bound. Let us check that (3.14) applies to this value of $m$.
@@
Thus, we may apply  (3.14). Choose a constant $\alpha\in(0,\frac{1}{4})$; in view of  (3.15), and (3.16), we get
@@ 3.2. Lower bounds for Kreiss conditions
Now, it is well-known that the uniform Kreiss constant of any operator is at most an (absolute) numerical constant times its absolute Cesáro constant (see e.g. the proofs of [Mont, Corollary 3.2] and [Quaestiones, Theorem 2]). In particular, we can find a constant $C>0$ only depending on $m$ such that $T_{0,c},T_{1,c},\dots,T_{p,c}$ and $\widetilde{T}_{p+1,c}$ all satisfy the uniform Kreiss condition with constant at most $C$, for any $0\leq c\leq 1.$ As (3.19) yields
@@
Using an analogous line of reasoning, it is possible to obtain a lower bound for $P_{S}$ as well. Indeed, we first recall that one of the key ingredients in the proofs of Theorems 1.2 and 3.3 is [[AbsolCesaromuller],Theorem 2.1], which asserts that the weighted backwards shift on $\ell^{2}$ with weights $w_{k}=\big(\frac{k}{k-1}\big)^{a},a>0,$ is absolutely Cesáro bounded. To treat the case of the strong Kreiss condition, one instead turns to [CohenCunyEisnerLin, Proposition 4.9]; this result shows that, if we choose $w_{k}=[\log(k+2)/\log(k+1)]^{a}$, then the associated weighted backwards shift satisfies the strong Kreiss condition, for any $a>0.$ Now, given $m\geq 1$ and $\varepsilon>0,$ one can define inductively, for any $c>0$ and $k>2$, $\widetilde{A}_{0,k}(c)=1$ and
@@
The proofs of Theorems 1.2 and 3.17 can then be repeated mutatis mutandis (with (2.2) replaced by the analogous inequality implied by [CohenCunyEisnerLin, Proposition 4.9]) to yield the existence of a constant $C_{S}=C_{S}(\varepsilon,m)$ such that
@@
This improves McCarthy’s growth rate of $\log^{1/2}(\log n)$ [McCarthy71lost, Example 2]. We omit the details.
@@
Now, by Paulsen’s theorem [Paulsensimilaritythm], it suffices to show that $\overline{\mathbb{D}}$ is a complete $K$-spectral set for $T$ for some $K>0.$ We will only show $K$-spectrality; the proof generalizes to the complete case by tensoring everything appropriately. Alternatively, we can apply a result by deLaubenfels [deLauben], which says that a Kreiss operator with finite peripheral spectrum is polynomially bounded if and only if it is completely polynomially bounded. (The same result for Ritt operators had been previously proved by Le Merdy [LMsimprob, Theorem 5.1]).
Continue the function $r(\cdot)$ to $\pi,\pi]$ by adding constant segments to its graph, so that it increases on $0,\pi]$, decreases on $\pi,0]$ and satisfies $r(-\pi)=r(\pi)=r(\delta_{+})=r(-\delta_{-})$. Put $1+\nu:=r(-\pi)=r(\pi)$ (so that $\nu>0$). By assumption, $\nu\leq\nu_{0}$. Extend $\gamma(\theta)=r(\theta)e^{-i\theta}$ to all $\theta\in[-\pi,\pi]$. Then, $\gamma$ is a Jordan curve that contains $\mathbb{D}$ in its interior and intersects $\mathbb{T}$ at $\{1\}$. Also, let $\{C_{n}\}_{n\geq 1}$ be a family of circles centered at the origin with radii $\{r_{n}\}_{n\geq 1}$ and such that
@@
To prove Lemma 4.1, we require the following well-known Poisson-type factorization. The proof is omitted (see e.g. [Crouzeix-Greenbaum, Lemma 7]).

### v2

@@ Abstract.
We also study operators that satisfy a variant of the above Kreiss condition where $K$ is replaced by $1+\varepsilon(|z|)$, where the positive continuous function $\varepsilon(|z|)$ tends to 0 as $|z|\to 1^{+}.$ We show that, if the spectrum of $T$ touches the unit circle only at a single point and the resolvent of $T$ satisfies a growth restriction along the unit circle, it is possible to choose $\varepsilon$ so that this Kreiss-type condition guarantees similarity to a contraction. At the core of our proof lies a positivity argument involving the double-layer potential operator. Counterexamples related to less restrictive choices of $\varepsilon$ are also provided.
@@ 1.1. The Kreiss condition
Kreiss [16] proved that every matrix that satisfies the above condition is power-bounded. More specifically, for every $N\times N$ matrix $A$ with Kreiss constant $K$, there exists $M(N,K)>0$ with
@@
In applications, however, the dimension $N$ may be very large (see e.g. [35]), so it is of interest to look for sharp constants $M(N,K)$.
Indeed, in the years after Kreiss published his result, various authors made successive improvements that reduced the value of $M(N,K)$ (see [34, Section 18] for the full history and references). Tadmor [33] was the first to show that $M(N,K)\leq CKN$ for some numerical constant $C>0$. Finally, Spijker [30], making crucial use of a Bernstein-type inequality for rational functions, proved that $M(N,K)\leq eKN$ and so, if $A$ is any $N\times N$ matrix with Kreiss constant $K,$ then
@@
LeVeque and Trefethen [18] showed that (1.2) is sharp in the sense that, for every $N\geq 1,$ there exists an $N\times N$ matrix $A_{N}$ with Kreiss constant $K_{N}$ such that
@@
By (1.2), we have $P(N,K)\leq eKN.$ What about lower bounds? The first results of this type are due to McCarthy-Schwartz [19] and van Dorsselaer-Kraaijevanger-Spijker [35], who established lower bounds for $P(N,K)$ that were logarithmic in $N$. Later, Spijker, Tracogna and Welfert [29] almost bridged the gap with (1.2) by showing that, given $\varepsilon>0,$ there exists $K=K(\varepsilon)>\pi+1$ and $C=C(\varepsilon)>0$ such that
@@
Their construction was subsequently refined by Nikolski [22], who showed that, for any $K>\pi+1$ and $N\geq 1,$
@@
More recently, the search for non-power-bounded operators that satisfy various Kreiss- and Cesàro-type conditions has focused on weighted shifts; see e.g. [1, 2, 5].
@@ 1.2. Small Kreiss constants
Unfortunately, all of the above-mentioned papers fall short of providing lower bounds for $P(N,K)$ for $K$ close to $1.$ In fact, it appears that the only time non-power-bounded operators with Kreiss-type constants close to $1$ were considered in the literature was in a paper of McCarthy [20] (see also the remark in [22, p. 374]), who proved that, for any $\varepsilon>0,$ there exists $C=C(\varepsilon)$ with
@@
    Let $\alpha\in(0,\frac{1}{4})$. There exist $M=M(\alpha)>0$ and $\varepsilon_{0}\in(0,1]$ such that, whenever $0<\varepsilon<\varepsilon_{0}$ and $\varepsilon\log N\geq M$, one has
@@ 1.3. The similarity problem
We now move from matrices to infinite-dimensional linear operators. Let $H$ be a complex Hilbert space and let $T\in\mathcal{B}(H)$ with $\sigma(T)\subset\overline{\mathbb{D}}$. The Kreiss condition
@@
(where $\nu>0$) no longer implies power-boundedness in this setting; classical counterexamples are contained in [9, 14]. On the other hand, if $C=1$, then $T$ is a $\rho$-contraction for some $\rho>1$, and so it is similar to a contraction (and thus power-bounded), no matter the value of $\nu$ (see e.g. [32]).
In the second half of our paper, we examine several conditions on $T$ that lie in-between the above $C=1$ and $C>1$ regimes, and determine whether they imply similarity to a contraction. We point out that similarity to a contraction (and, more generally, $K$-spectral estimates) has already been linked to resolvent conditions several times in the literature, see e.g. [3, 4, 7, 8, 10]. For more references and recent developments concerning the similarity problem, especially in relation to operator semigroups, see [23]. One of the conditions we are interested in, naturally motivated by the consideration of “small” Kreiss constants, is
@@
We also consider operators $T$ such that $\lim_{n\to\infty}\|T^{n}\|=1$. Notice that if $\|T^{m}\|\leq 1$ for some $m\in\mathbb{N}$, then $T$ is similar to a contraction, see  [[15], discussion of Problem 6]. It is easy to see that the condition $\lim_{n\to\infty}\|T^{n}\|=1$ implies (1.9) for a certain function $\varepsilon(x)$ vanishing at $0^{+}$. By adapting Pisier’s example of an operator that is polynomially bounded but not similar to a contraction [27], we show
@@
Let $T\in\mathcal{B}(H)$ with $\sigma(T)\subset\mathbb{D}\cup\{1\}$, with $\gamma:[-\delta_{-},\delta_{+}]\to\mathbb{C}$ a v-type curve. Assume that $T$ satisfies the Kreiss condition and, moreover, $\varepsilon:(0,\nu]\to(0,\infty)$ is a continuous increasing function such that $\varepsilon(0^{+})=0$ and the resolvent estimate (1.9) holds for all $\lambda$ in the set determined by $\gamma$ (excluding $\lambda=1$). Set $r(\theta)=|\gamma(\theta)|$. If
@@
for all $\theta$ (see e.g. [28, Section 1.2]). Thus, in view of Theorem 1.5, we immediately obtain
@@
We note that Ritt operators, while power-bounded, are not, in general, similar to contractions [17, Proposition 5.2].
@@ 1.4. Outline
Our paper is organized as follows: in Section 2, we briefly go over the definitions and the relations between those Kreiss- and Cesáro-type conditions that are relevant to this work. We also record an inequality from [1] that will be very useful in the proof of Theorem 1.2. Section 3 contains the proofs of Theorem 1.2 and Theorem 3.3, and, in addition, a lower bound corresponding to the strong Kreiss condition (Remark 3.7). Theorem 1.5 is proved in Section 4. Section 5 contains the proofs of Theorems 5.1 and 1.3. Section 6 concludes with some open questions related to our investigations.
@@ 2. Kreiss- and Cesàro-type conditions
Clearly, power-boundedness implies all of the above conditions. Absolute Cesàro boundedness and the strong Kreiss condition are both strictly stronger than the uniform Kreiss condition, see [1, 12, 21], while the two are independent of one another [5]. Further, the uniform Kreiss condition is equivalent [21] to the existence of a constant $C>0$ such that, for all $n\geq 1$ and $|\lambda|=1,$
@@
and is thus stronger than Cesáro boundedness (and also, trivially, stronger than the Kreiss condition). The Kreiss condition and Cesàro boundedness are independent [31]. Finally, while the strong Kreiss condition does not imply power-boundedness [5], we observe that the similar-looking condition
@@
is, in fact, equivalent to power-boundedness [11, 24].
@@
An interesting example of an operator that is absolutely Cesáro but not power bounded is given in [[1], Theorem 2.1] (we only record the $\ell^{2}$-version here):
@@ 3.1. Lower bounds for the absolute Cesàro condition
$P_{AC}(N^{\prime\prime},1+\varepsilon)\geq 2^{-m}C(\varepsilon,m)\log^{m}(N^{\prime\prime}),$ (3.14)
for all $N^{\prime\prime}\geq m^{2}2^{2m}$. Indeed, let $N^{\prime\prime}$ satisfy $N^{\prime\prime}\geq m^{2}2^{2m},$ with $N^{\prime}$ the largest integer such that $N^{\prime}\leq N^{\prime\prime}$ and $N^{\prime}$ is a multiple of $2^{m}$. Since $P_{AC}(N^{\prime\prime},K)$ is increasing with respect to $N^{\prime\prime}$ (assuming $K$ fixed), (3.13) allows us to write
$P_{AC}(N^{\prime\prime},1+\varepsilon)$ $\geq P_{AC}(N^{\prime},1+\varepsilon)$
@@
$\geq C(\varepsilon,m)\log^{m}(N^{\prime\prime}/2)$
$\geq 2^{-m}C(\varepsilon,m)\log^{m}(N^{\prime\prime}),$
@@
Set $m=\lfloor\frac{v}{2\log v}\rfloor$, where $v=\log(\varepsilon\log N)$. This choice of $m$ is, in a sense, close to the value that maximizes this last lower bound. Let us check that $N\geq m^{2}2^{2m}$, which will allow us to apply (3.14).
@@
Thus, we may apply  (3.14). Choose a constant $\alpha\in(0,\frac{1}{4})$; in view of  (3.15) and (3.16), we get
@@ 3.2. Lower bounds for Kreiss conditions
Now, it is well-known that the uniform Kreiss constant of any operator is at most an (absolute) numerical constant times its absolute Cesáro constant (see e.g. the proofs of [21, Corollary 3.2] and [13, Theorem 2]). In particular, we can find a constant $C>0$ only depending on $m$ such that $T_{0,c},T_{1,c},\dots,T_{p,c}$ and $\widetilde{T}_{p+1,c}$ all satisfy the uniform Kreiss condition with constant at most $C$, for any $0\leq c\leq 1.$ As (3.19) yields
@@
Using an analogous line of reasoning, it is possible to obtain a lower bound for $P_{S}$ as well. Indeed, we first recall that one of the key ingredients in the proofs of Theorems 1.2 and 3.3 is [[1],Theorem 2.1], which asserts that the weighted backwards shift on $\ell^{2}$ with weights $w_{k}=\big(\frac{k}{k-1}\big)^{a},a>0,$ is absolutely Cesáro bounded. To treat the case of the strong Kreiss condition, one instead turns to [5, Proposition 4.9]; this result shows that, if we choose $w_{k}=[\log(k+2)/\log(k+1)]^{a}$, then the associated weighted backwards shift satisfies the strong Kreiss condition, for any $a>0.$ Now, given $m\geq 1$ and $\varepsilon>0,$ one can define inductively, for any $c>0$ and $k>2$, $\widetilde{A}_{0,k}(c)=1$ and
@@
The proofs of Theorems 1.2 and 3.17 can then be repeated mutatis mutandis (with (2.2) replaced by the analogous inequality implied by [5, Proposition 4.9]) to yield the existence of a constant $C_{S}=C_{S}(\varepsilon,m)$ such that
@@
This improves McCarthy’s growth rate of $\log^{1/2}(\log n)$ [20, Example 2]. We omit the details.
@@
Now, by Paulsen’s theorem [25], it suffices to show that $\overline{\mathbb{D}}$ is a complete $K$-spectral set for $T$ for some $K>0.$ We will only show $K$-spectrality; the proof generalizes to the complete case by tensoring everything appropriately. Alternatively, we can apply a result by deLaubenfels [7], which says that a Kreiss operator with finite peripheral spectrum is polynomially bounded if and only if it is completely polynomially bounded. (The same result for Ritt operators had been previously proved by Le Merdy [17, Theorem 5.1]).
Continue the function $r(\cdot)$ to $\pi,\pi]$ by adding constant segments to its graph, so that it increases on $0,\pi]$, decreases on $\pi,0]$ and satisfies $r(-\pi)=r(\pi)=r(\delta_{+})=r(-\delta_{-})$. Without loss of generality, we may assume that $\varepsilon$ is defined on the interval $(0,\nu]$, where $1+\nu=r(-\pi)=r(\pi)$ (note $\nu>0$). Extend $\gamma(\theta)=r(\theta)e^{-i\theta}$ to all $\theta\in[-\pi,\pi]$. Then, $\gamma$ is a Jordan curve that contains $\mathbb{D}$ in its interior and intersects $\mathbb{T}$ at $\{1\}$. Also, let $\{C_{n}\}_{n\geq 1}$ be a family of circles centered at the origin with radii $\{r_{n}\}_{n\geq 1}$ and such that
@@
To prove Lemma 4.1, we require the following well-known Poisson-type factorization. The proof is omitted (see e.g. [6, Lemma 7]).